OpenStudy (anonymous):

p.t 7/3 root5 is an irrational no

OpenStudy (anonymous):

step by step i need d answer cn any1 plss help me .................. :(

OpenStudy (amorfide):

 is this your question?

Parth (parthkohli):

I assume that "p.t" means "prove that"?

OpenStudy (anonymous):

yaaaaaaaaaa @ParthKohli and @amorfide

OpenStudy (amorfide):

I don't know the proof sorry

OpenStudy (anonymous):

ok @ParthKohli cn u do it

OpenStudy (anonymous):

@Hero cn u help mee

Parth (parthkohli):

Hmm. Assume that your number is rational.\[{7 \over 3\sqrt5} = {p \over q} \,\wedge \gcd (p,q) = 1\]

Parth (parthkohli):

\[{49 \over45 } = {p^2 \over q^2}\]\[49q^2 =45p^2 \]This doesn't seem like the conventional proof. :(

OpenStudy (anonymous):

is it correct

Parth (parthkohli):

No, I need to find a contradiction which I am unable to do.

OpenStudy (anonymous):

uff will u try 2 doo ma frnddd

OpenStudy (anonymous):

7/3sqrt5 = p/q -> sqrt 5 = 7q/3p and square -> 5=49q^2/9p^2 so p^2/q^2 = 49/45 but no integer squares to 45

Parth (parthkohli):

Oh, so that was the contradiction!